OpenStudy (anonymous):

I am trying to understand what the purpose/application of factoring as in quadratics is.

OpenStudy (anonymous):

There are several reasons why you would factor an quadratic equation, however, the first that comes to my mind is solving for roots or for the x-value(s). Ex) x^2-2x=-1 x^2-2x+1=0 If we factor: (x-1)(x-1)=0 From this we can see that x=1, so that's our x-value for this equation. You can double check that by plugging back in: 1-2+1=0 ... Check